In a rectangle, the ratio of the length to the width is 5:4. If the rectangle is 30 inches long, find its width. (Points: 1)
Answer by edjones(8007)

Let x=width
x/30=4/5 Proportion.
30x/30=30*4/5
x=24
.
Ed